OpenStudy (anonymous):

What is the diameter of a cone with height 8 m and volume 150m3 ?

OpenStudy (radar):

Formula for volume of a cone. V = (pi r^2 h)/3 150 =(pi r^2 h)/3 150 =(pi r^2 8)/3 solve for r then 2r = diameter.

OpenStudy (dayianaforever):

answer 15

OpenStudy (radar):

I got 8.463 meters 150 =(pi r^2 8)/3 multiplying by 3 450 = 8 pi r^2 dividing by 8 pi r^2 =17.905 taking square root of both side r = 4.23142 diameter is times 2 D= 8.463 meters